OpenStudy (anonymous):

Solve the equation. 3(x – 9) = 30 x = 1 x = 13 x = 19 x = 57

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) for 3(x - 9) = 30 first distribute so 3*x - 3*9 = 30 3x - 27 = 30 then add 27 to both sides +27 +27 3x = 57 then divide both sides by 3 so x = 57/3 = ?

OpenStudy (anonymous):

@jigglypuff314 does that mean divided 53 by 3 ? if it does then that equals 17.6?

jigglypuff314 (jigglypuff314):

57 divided by 3 :)

jigglypuff314 (jigglypuff314):

becaues when you add 27 to both sides 30 + 27 = 57

OpenStudy (anonymous):

oh sorry thats 19 then

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

so x is 19

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

thanks ! gold medal to you !!

jigglypuff314 (jigglypuff314):

glad I could help :)